OpenStudy (anonymous):

ln (x-4) - ln 9=3 Please explain how to solve it too because I can't remember anything from this section!!! The answer is supposed to be 9e^3+4 Thank you!

Parth (parthkohli):

Hi, can you simplify the left-hand side using\[\ln(a) - \ln(b) = \ln\left(\dfrac{a}{b}\right)\]

Parth (parthkohli):

Also, you can write the right-hand side as \(\ln(e^3)\). Are you getting it?

OpenStudy (anonymous):

I think I kind of understand it now. I'll just look in the book's examples to see if I get how it works :) Thank you!

Parth (parthkohli):

Can you solve it now? I hope you can :-|